OpenStudy (anonymous):

evaluate the function f(x)=x^2+4x-2;f(-3)

OpenStudy (anonymous):

(-3)^2+4(-3)-2=9-12-2=-5 The answer is -5.

OpenStudy (saifoo.khan):

-5

OpenStudy (anonymous):

thanks for the help